OpenStudy (anonymous):

Hey everyone, Could someone help me on this question , i got stumped again =/ Simplify: [2(x+1)^3/2 (6x+4) - 3(3x^2+4x)(x+1)^1/2]/[4(x+1)^3] Answer: 3x^2+8x+8/ 4(x+1)^5/2

OpenStudy (anonymous):

\[2(x+1)^{3/2}(6x+4) -3(3x ^{2}+4x)(x+1)^{1/2} / 4(x+1)^{3}\]

OpenStudy (anonymous):

your going to simplify it?

OpenStudy (anonymous):

Yea, kinda got stumped on how they got 5/2 as the exponent on the bottom

OpenStudy (anonymous):

Well, lets just consider that part \[\frac{(x+1)^{1/2}}{(x+1)^3}=(x+1)^{1/2-3} = (x+1)^{-5/2}\] Hmmm, so your wondering why they get a - there? Because that's what I get :)

OpenStudy (anonymous):

Oh ic, though the answer is +5/2 as the exponent, wierd.

OpenStudy (anonymous):

weird*

OpenStudy (anonymous):

It's correct, \[(x+1)^{-5/2} = \frac{1}{(x+1)^{5/2} }\]

OpenStudy (anonymous):

Thought it was a product you had written on top.

OpenStudy (anonymous):

Ohh okay then that makes sense thanks a lot for your help

OpenStudy (anonymous):

no worries:)